• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name T. Badman

    Log entry time 14:23:32 on March31,2012

    Entry number 368648

    keyword=late- MCC finished tuning beam

    MCC finished tuning beam and noticed 1H01 was 3mm off from the setting before moller, all the beam position values have been restored. The harp scan was unsuccessful because BPMA and BPMB saturated. MCC would like to adjust the BPM gains at some point.